For the demand equation P = 36 - 2Q, what Price will maximize total revenue?
Answer #1

Answer

Total revenue is maximum when marginal revenue is zero

Total revenue =P*Q=36Q-2Q^2

change in total revenue is marginal revenue and the change in function found by differentiation

MR=dTR/dQ=36-4Q

equating to zero

36-4Q=0

4Q=36

Q=9

P=36-2*9

P=$18

the revenue is maximized when P=$18